# The Birth of Be Star Disks II. A High-Resolution Spectroscopic Campaign and TESS Observations of an Outburst of the Classical Be star Pavonis

### v1

@@ 5.2 High-frequency stellar variability
The character and details of the LPVs differ from one night to the next.
@@
There are no high-frequency signals detected in TESS.
@@
The features are in general highly structured.
@@
Interpreting the physical cause of these variations is not entirely straightforward, as there are several distinct processes that can cause line profile variations. A star with an inhomogeneous surface (e.g. with temperature and/or chemical spots) can exhibit LPVs with a frequency equal to the stellar rotation (or over a range of frequencies if the star has surface differential rotation). Such spots may vary over time (appearing, migrating, and disappearing as with Sun spots), or they may be apparently constant over long time baselines. Variations in the circumstellar environment can cause LPVs, too, and these variations can be intrinsic or related to projection effects from obscuration of (parts of) the observed stellar disk and/or self-obscuration as material rotates or orbits around the star. Pulsation is another mechanism, as there are several types of pulsation with many different geometric possibilities which can induce LPVs with a wide range of patterns and timescales.
It is useful to keep in mind that a typical set of continuous observations in our data is 5 hours long, which is approximately 10% of the rotation period, or 20% of the near-surface orbital period (both of these estimated from the parameters of 2016A&A...595A.132Z).
@@
In Si III 5739, which is the only line apparently unaffected by circumstellar material, there do not seem to be any major differences in the appearance of the fast-moving features between sets A – D. All variations are confined to within km s-1, and a combination of prograde and retrograde features are seen. These LPVs bear resemblance to high-order (in terms of ) pulsation.
@@
It is not uncommon for early-type Be stars to exhibit high-frequency p modes (i.e. Cephei pulsation). The characteristics of the fast LPVs in Pav seem different for the following reasons. For example, the early-type Be stars Aqr and BZ Cru both exhibit high-frequency p modes, which i) are detected in photometry, and ii) appear as regular periodic patterns in line profiles (2020MNRAS.494..958N). In Pav, there are no corresponding photometric frequencies (despite the high precision of TESS, although geometric cancellation can make high-order modes all but invisible to photometry), and the LPV patterns differ significantly from night to night. For instance, on the night of 2023 July 28, four or five features move roughly in parallel from blue to red, but possibly with changing acceleration throughout the night. On 2023 Aug 12, there are only two or three features, primarily only on the red half of the line, which hardly move across the line profile but rather appear and disappear at near-constant velocity. This could represent a crossing of prograde and retrograde moving features at that time. On 2023 July 22, there are again four or five features, but moving with clearly different accelerations, most being prograde, but some being retrograde. The work of 2016A&A...595A.132Z finds that Pavonis has an inclination of 36∘ 10∘ and a of 145 km s-1 12 km s-1. Considering the low inclination angle, the retrograde moving features could actually be propagating in the direction of stellar rotation, but are visible from the “back” side of the star and thus move from red to blue in the observed line profiles.
@@
The TESS observations show no indication of photometric signals associated with these fast LPVs, indicative that these should be higher-order LPVs. As a first attempt, we modeled several different combinations of pulsation geometries, and found the closest qualitative match was with an = 7, = -7 sectoral prograde mode in BRUCE/KYLIE (2014ascl.soft12005T) as the observed frequency is coursely proportional to . The resulting variations qualitatively match the dominant features in Si III 5739 on the night of 2023 July 22 (Fig. 6). However, this single mode of course cannot explain the night-to-night differences or the sometimes retrograde or near-stationary features. Like with the slow LPVs, a dedicated effort to model these features is needed to understand their details and is beyond the scope of this paper. It is possible that multiple coherent modes are present at the same time, constructively and destructively interfering in ways that at first glance appear random, but could be understood if the constituent mode frequencies and geometries are known. Alternatively, the implied high-order pulsation could be non-coherent. It is noteworthy that these fast-moving relatively sharp features do not seem to appear in any of the spectra presented in 2011A&A...533A..75L. We do not doubt the reality of these signals in the 2023 observing run, since they are clearly seen in data from two different instruments and in every long observing sequences.
@@
Qualitatively similar features have been observed in a few other early-type Be stars, including Cas and Tau (1988PASP..100..233Y; 1990ApJS...74..595Y; 1999ApJ...517..866S; 1998ApJ...507..945S). In both of these cases, the stars possessed strong disks at the time of observations, so that both stellar and circumstellar interpretations were valid. However, for Pav, it is safe to attribute these to pulsational features on the stellar surface thanks to observations of the purely photospheric Si III 5739 line and the absence of any circumstellar material prior to the outburst.
@@
During the rising phase of the outburst (2023 July 27, 28 in Fig. 6), the H I and He I lines, which are sensitive to circumstellar material, show clear variations well outside of the stellar sin, especially evident on 2023 July 27 on the red side of the line. As discussed in Sec. 2, the circumstellar material is still inhomogeneously distributed at these early times prior to circularization. While the underlying pulsational variability makes it difficult to track the motion of any gas during these several hour sequences, the super- features (which are not observed in the photospheric Si III 5739 line) can safely be attributed to circumstellar material. While this would nominally violate the conservation of angular momentum, in these cases, the material is conserving angular momentum by orbiting the star. Recent simulations of inhomogeneous mass ejection leading to disk build-up in Be stars show that prior to circularization, and while mass ejection is ongoing, the gas dynamics are complex and do not resemble circular Keplerian orbits (Rubio et al., priv. comm.). High-cadence spectroscopic observations during the early stages of mass ejection such as for Pav will be important in refining models of this sort to better understand the initial conditions of ejected material, hopefully elucidating the stellar processes that give rise to the ejection of mass.
5.3 Circumstellar environment
@@
During our intensive spectroscopic observing campaign, and overlapping with space photometry from TESS in June and July 2023, the Pav system transitioned from a disk-less state by means of a discrete mass ejection episode and the formation of a new disk. The disk build-up phase lasted for only around 3–5 days for all lines sensitive to emission (Fig. 2 and the exponential fit in Table 2), during which time the emission levels of H I and He I lines increased while emission asymmetry oscillated with a period of about 1.25 d ( d-1; Fig. 3, Table 2). After peak emission levels were reached, the dissipation proceeded more quickly for He I lines compared to H I. About 30–40 days after the start of the event, emission features were no longer present in any lines at optical wavelengths. Only the first few days of the event were captured by TESS, so the timescale of the build-up and dissipation phases in continuum flux remain unknown. Overall, the behavior related to the short outburst in Pav is much the same as in comparable events in other Be stars (Paper 1).
@@
The observed behavior of the emission levels of the H I and He I lines is consistent with the generally accepted picture of Be disk build-up and dissipation. During the build-up phase, the inner volume of the disk fills up relatively quickly. At these early times, all observables originate from near to the star (insufficient time has elapsed for material to be viscously transported to large distances), and we observe the same build-up times in both the H I and He I emission. When mass ejection turns off, the inner-most regions of the disk dissipate more quickly with the bulk of the recently ejected material falling back on to the star (e.g. 2021ApJ...912...76M; 2021ApJ...909..149G). Helium emission tends to arise from hot dense material relatively close to the star, so He emission levels drop relatively quickly during the dissipation phase. Furthermore, we see that the higher-energy He I at 4713 and 4921 Å dissipate quicker than the lower-energy He I 5876 and 6678 lines. Hydrogen emission, on the other hand, can arise from cooler and less dense regions of the disk farther from the star. All relevant time dependent events, in particular the rate of change in density caused by variations in the stellar mass ejection rate, proceed more slowly at larger radial distances (2012ApJ...756..156H). Thus, the slower drop in H I emission, relative to He I, is consistent with the picture of Be disks dissipating from the ‘inside-out’. In this picture, the density of the inner disk drops quickly, while the decreasing density in more outer parts of the disk lags behind. In fact, this behavior can be directly observed for H in Fig. 1. The high-velocity wings of H, which originate in the more inner disk region, dissipate more quickly than the emission closer to the line center which preferentially forms at larger radii in the disk where Keplerian orbital speeds are lower.
6 Conclusions
@@
The findings in this paper have detailed pulsational properties both photometrically and spectroscopically at times when a disk was built and dissipated around the classical Be star Pav. The data set presented will provide theorists a rich landscape with which to model how small ejections enter the Be star’s disk and then dissipate with time. Furthermore, the pulsational properties of the star will be of utmost importance to understand this star’s properties and how they relate to the disk evolution. Lastly, some prevailing thoughts for Be stars may need to be reconsidered given the way in which the fast line profile variations are similar at the times before, during, and after the outburst. Our main findings are as follows:
@@
The timescale for disk build-up was consistent in all lines that were analyzed (4 days; Sec. 2). After reaching peak emission levels, the decay proceeded more quickly for He I lines (4 days) compared to H (6 days) and H (9 days).
@@
During the first several days of the outburst, the emission asymmetry oscillated with a period of 1.25 days in all measured lines with emission features (ec. 2). After 10 days from the start of the event, emission features had settled into a symmetric profile.
@@
Besides the outburst, a signal with a 6.13 day period is the most prominent variability in the TESS photometry. A spectroscopic signal at this period was noted in 2011A&A...533A..75L, and is also detected in the LPVs of our spectra. The frequency of this signal is consistent with the difference between two higher-frequency signals (at 1.644 d-1 and 1.485 d-1). The variations associated with this difference frequency are remarkably stable, apparently over many years (Sec. 5.1), and also before, during, and after the outburst.
@@
Fast-moving, highly structured, and apparently changing features were evident in line profiles in all nine nights where Pav was observed for several consecutive hours (Sec. 5.2). These appear as a combination of prograde and retrograde features, were seen before, during, and after the outburst, and have no photometric counterpart.
@@
2011A&A...533A..75L found four frequencies in their spectroscopic analysis of Pav. We confirm that three of these are present in the TESS photometry.

### v2

@@ V.2 High-frequency stellar variability
The character and details of the LPVs differ from one night to the next. It may be possible to model these as several periodic LPVs, but that will likely require detailed modeling and further observations to determine the timescale for each component.
@@
There are no high-frequency signals ( 2 d-1) detected in TESS.
@@
The LPVs are in general highly structured.
@@
Interpreting the physical cause of these variations is not entirely straightforward, as there are several distinct processes that can cause line profile variations. A star with an inhomogeneous surface (e.g. with temperature and/or chemical spots) can exhibit LPVs with a frequency equal to the stellar rotation (or over a range of frequencies if the star has surface differential rotation) (Semenova, 2006). Such spots may vary over time (appearing, migrating, and disappearing as with Sun spots), or they may be apparently constant over long time baselines. Variations in the circumstellar environment can cause LPVs, too, and these variations can be intrinsic or related to projection effects from obscuration of (parts of) the observed stellar disk and/or self-obscuration as material rotates or orbits around the star. Pulsation is another mechanism, as there are several types of pulsation with many different geometric possibilities which can induce LPVs with a wide range of patterns and timescales (Telting, 2003).
It is useful to keep in mind that a typical set of continuous observations in our data is 5 hours long, which is approximately 10% of the rotation period, or 20% of the near-surface orbital period (both of these estimated from the parameters of Zorec et al., 2016).
@@
In Si III 5739, which is the only line apparently unaffected by circumstellar material, there do not seem to be any major differences in the appearance of the fast-moving features between sets A – D. All variations are confined to within km s-1, and a combination of prograde and retrograde features is seen. These LPVs bear resemblance to high-order (in terms of ) pulsation (Reid, 1994).
@@
It is not uncommon for early-type Be stars to exhibit high-frequency p-modes (i.e. Cephei pulsation with frequencies between and cycles d-1 Stankov and Handler, 2005). The characteristics of the fast LPVs in Pav seem different for the following reasons. For example, the early-type Be stars Aqr and BZ Cru both exhibit high-frequency p modes, which i) are detected in photometry, and ii) appear as regular periodic patterns in line profiles (Nazé et al., 2020). In Pav, there are no corresponding photometric frequencies in Fig. 4 (despite the high precision of TESS, although geometric cancellation can make high-order modes all but invisible to photometry), and the LPV patterns differ significantly from night to night, as expected for multi-mode pulsations. For instance, on the night of 2023 July 28 (row 7 in Fig. 6), four or five features move roughly in parallel from blue to red, but possibly with changing acceleration throughout the night. On 2023 Aug 12 (row 8 of Fig. 6), there are only two or three features, primarily only on the red half of the line, which hardly move across the line profile but rather appear and disappear at near-constant velocity. This could represent a crossing of prograde and retrograde moving features at that time. On 2023 July 22, (row 4 of Fig. 6), there are again four or five features, but moving with clearly different accelerations, most being prograde, but some being retrograde. The work of Zorec et al. (2016) finds that Pavonis has an inclination of 36∘ 10∘ and a of 145 km s-1 12 km s-1. Considering the low inclination angle, the retrograde moving features could actually be propagating in the direction of stellar rotation, but are visible from the “back” side of the star and thus move from red to blue in the observed line profiles.
@@
The TESS observations show no indication of photometric signals associated with these fast LPVs, indicative that these should be higher-order LPVs. As a first attempt, we modeled several different combinations of pulsation geometries, and found the closest qualitative match was with an = 7, = -7 sectoral prograde mode in BRUCE/KYLIE (Townsend, 2014) as the observed frequency is coursely proportional to . The resulting variations qualitatively match the dominant features in Si III 5739 on the night of 2023 July 22 (Fig. 6, column 4, row 4). However, this single mode of course cannot explain the night-to-night differences or the sometimes retrograde or near-stationary features. Like with the slow LPVs, a dedicated effort to model these features is needed to understand their details and is beyond the scope of this paper. It is possible that multiple coherent modes are present at the same time, constructively and destructively interfering in ways that at first glance appear random, but could be understood if the constituent mode frequencies and geometries are known. Alternatively, the implied high-order pulsation could be non-coherent. It is noteworthy that these fast-moving relatively sharp features do not seem to appear in any of the spectra presented in Levenhagen et al. (2011). We do not doubt the reality of these signals in the 2023 observing run, since they are clearly seen with both NRES and CHIRON in every long observing sequence.
@@
Qualitatively similar features have been observed in a few other early-type Be stars, including Cas and Tau (Yang et al., 1988, 1990; Smith and Robinson, 1999; Smith et al., 1998). In both of these cases, the stars possessed strong disks at the time of observations, so that both stellar and circumstellar interpretations could be valid. However, for Pav, it is safe to attribute these to pulsational features on the stellar surface thanks to observations of the purely photospheric Si III 5739 line and the absence of any circumstellar material prior to the outburst.
@@
During the rising phase of the outburst (2023 July 27, 28, rows 6 and 7 in Fig. 6), the H I and He I lines, which are sensitive to circumstellar material, show clear variations well outside of the stellar sin, especially evident on 2023 July 27 on the red side of the line. As discussed in Sec. II, the circumstellar material is still inhomogeneously distributed at these early times prior to circularization. While the underlying pulsational variability makes it difficult to track the motion of any gas during these several hour sequences, the super- features (which are not observed in the photospheric Si III 5739 line) can safely be attributed to circumstellar material. While this would nominally violate the conservation of angular momentum, in these cases, the material is conserving angular momentum by orbiting the star by being viscously advected to larger distances. Recent simulations of inhomogeneous mass ejection leading to disk build-up in Be stars show that prior to circularization, and while mass ejection is ongoing, the gas dynamics are complex and do not resemble circular Keplerian orbits (Rubio et al., priv. comm.). High-cadence spectroscopic observations during the early stages of mass ejection such as for Pav will be important in refining models of this sort to better understand the initial conditions of ejected material, hopefully elucidating the stellar processes that give rise to the ejection of mass.
V.3 Circumstellar environment
@@
During our intensive spectroscopic observing campaign, and overlapping with space photometry from TESS in June and July 2023, the Pav system transitioned from a disk-less state by means of a discrete mass ejection episode to the formation of a new disk. The disk build-up phase lasted for only around 3–5 days for all lines sensitive to emission (Fig. 2 and the exponential fit in Table 2), during which time the emission levels of H I and He I lines increased while emission asymmetry oscillated with a period of about 1.25 d ( d-1; Fig. 3, Table 2). After peak emission levels were reached, the dissipation proceeded more quickly for He I lines compared to H I. About 30–40 days after the start of the event, emission features were no longer present in any lines at optical wavelengths. Only the first few days of the event were captured by TESS, so the timescale of the build-up and dissipation phases in continuum flux remain unknown. Overall, the behavior related to the short outburst in Pav is much the same as in comparable events in other Be stars (Paper I).
@@
The observed behavior of the emission levels of the H I and He I lines is consistent with the generally accepted picture of Be disk build-up and dissipation. During the build-up phase, the inner volume of the disk fills up relatively quickly. At these early times, all observables originate from near to the star (insufficient time has elapsed for material to be viscously transported to large distances), and we observe the same build-up times in both the H I and He I emission. When mass ejection turns off, the inner-most regions of the disk dissipate more quickly with the bulk of the recently ejected material falling back on to the star (e.g., Marr et al., 2021; Ghoreyshi et al., 2021). Helium emission tends to arise from hot dense material relatively close to the star, so He emission levels drop relatively quickly during the dissipation phase. Furthermore, we see that the higher-excitation He I at 4713 and 4921 Å dissipate quicker than the lower-excitation He I 5876 and 6678 lines. Hydrogen emission, on the other hand, can arise from cooler and less dense regions of the disk farther from the star. All relevant time dependent events, in particular the rate of change in density caused by variations in the stellar mass ejection rate, proceed more slowly at larger radial distances (Haubois et al., 2012). Thus, the slower drop in H I emission, relative to He I, is consistent with the picture of Be disks dissipating ‘inside-out’. In this picture, the density of the inner disk drops quickly, while the decreasing density in more outer parts of the disk lags behind. In fact, this behavior can be directly observed for H in Fig. 1. The high-velocity wings of H, which originate in the more inner disk region, dissipate more quickly than the emission closer to the line center which preferentially forms at larger radii in the disk where Keplerian orbital speeds are lower.
VI Conclusions
@@
The findings in this paper have detailed pulsational properties both photometrically and spectroscopically at times when a disk was built and dissipated around the classical Be star Pav. The data set presented will provide theorists a rich landscape with which to model how small ejections enter the Be star’s disk and then dissipate with time. Furthermore, the pulsational properties of the star will be important to understand this star’s properties and how they relate to the disk evolution. Lastly, some prevailing thoughts for Be stars may need to be reconsidered given the way in which the fast line profile variations are not similar at the times before, during, and after the outburst. Our main findings are as follows:
@@
The timescale for disk build-up was consistent in all lines that were analyzed (4 days; Sec. II). After reaching peak emission levels, the decay proceeded more quickly for He I lines (4 days) compared to H (6 days) and H (9 days).
@@
During the first several days of the outburst, the emission asymmetry oscillated with a period of 1.25 days in all measured lines with emission features (Sec. II), which we interpret as orbiting circumstellar material. After 10 days from the start of the event, emission features had settled into a symmetric profile.
@@
Besides the outburst, a signal with a 6.13 day period is the most prominent variability in the TESS photometry. A spectroscopic signal at this period was noted in Levenhagen et al. (2011), and is also detected in the LPVs of our spectra. The frequency of this signal is consistent with the difference between two higher-frequency signals (at 1.644 d-1 and 1.485 d-1), which could be NRPs. The variations associated with this difference frequency are remarkably stable, apparently over many years (Sec. V.1), and also, spectroscopically, before, during, and after the outburst.
@@
Fast-moving, highly structured, and apparently changing features were evident in line profiles in all nine nights where Pav was observed for several consecutive hours (Sec. V.2). These appear as a combination of prograde and retrograde features, were seen before, during, and after the outburst, and have no photometric counterpart.
@@
Levenhagen et al. (2011) found four frequencies in their spectroscopic analysis of Pav. We confirm that three of these (0.163 d-1, 0.82 d-1, and 1.644 d-1) are present in the TESS photometry.
All authors contributed to this writing. SSN led the analysis and measurements and prepared most of the manuscript under the guidance and mentorship of NDR. JLB provided guidance in the analysis and data, also scheduling the LCO observations and leading the analysis of the TESS light curve. NDR scheduled and planned the CHIRON observations. SGF provided the foundation of the MCMC code used in this manuscript.